# Assessing the viability of non-light water reactor concepts for electricity and heat generation in decarbonized energy systems

### v1

@@ 2.2.2 Reactors with Fast Neutron Spectra
Lead-cooled fast reactors (LFR) operate similarly to SFRs, but use lead as a coolant. Lead is less reactive than sodium and could thus reduce operational risks and necessary safety barriers [64]. However, operational experience is almost exclusively limited to Russia, where LFRs were experimented with for submarine propulsion, and a single reactor is currently under construction [60]. Further, a demonstration LFR with an accelerator-driven system is currently under development in Belgium (MYRRHA) [63]. Due to limited operational experience, cost information is scarce. However, it is expected that LFRs could be cheaper than SFRs due to reduced safety risks, and if a closed fuel cycle can be implemented, LFRs might also compete with current LWRs [64].
@@ 2.2.4 Molten salt reactors
Molten salt reactors (MSR) use molten salt as a coolant [68]. MSRs were initially developed for aircraft propulsion, and concepts vary regarding moderator, fuel, neutron spectrum, and fissile material. There are no commercially operational MSRs, but several concepts are currently under development in China, Switzerland, and Russia, amongst others [63]. MSRs could become applicable for heat provision in decarbonized energy systems as the proposed salts melt at around 500°C and boil at around 1400°C, allowing for outlet temperatures ranging from 600 to 700°C [68, 63]. Most MSR concepts are still in the early development stages, so information on their economic competitiveness is limited. While in the 1970s, capital costs of MSR were estimated in the same range as LWRs [69], current concepts are estimated to be cheaper. For example, the so-called "Advanced High-Temperature Reactor," also referred to as "Fluoride High-Temperature Reactor," is estimated to cost less than 1000 USD per kW, and the ThorCon concept MSR, developed by company Martingale, is envisioned to produce electricity for as little as 30 to 50 USD per MWh [70, 63].
2.3 Nuclear Cost Data Projections
@@
Over the last decade, nuclear power has gained momentum in energy scenarios as a low-carbon technology. For example, over 80% of scenarios collected in the 6th Assessment Report of the IPCC show that to limit global warming to 1.5°C by the year 2100, it is necessary to increase absolute electricity generation from nuclear with an average annual growth rate of 3% [1, 7]. These models assume substantial overnight construction cost reductions for nuclear in the coming decades down to 5,000 or even 2,000 USD per kW [20] and highlight the necessity of cost reductions for nuclear to be competitive in energy systems with high shares of renewables [2, 71].
@@
Thus, to assess the economic efficiency of non-light water reactors in the energy system model (see Section 3.1), it is first necessary to determine relevant cost data and other relevant information on potentially viable reactor types. Consequently, we conducted a literature analysis of cost projections and assumptions for various reactor types. Previous research has shown that for high-capacity LWRs, cost projections and experience differ strongly [8]. For SMR concepts, Steigerwald et al. [5] show that to-be-expected costs and industry expectations also vary substantially.
@@
The here presented cost analysis is based on 51 references [72, 73, 3, 63, 74, 75, 76, 77, 78, 66, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 16, 94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105, 106, 67, 15, 107, 108, 109]. About 60% are so-called grey literature, mostly industry and research organization reports. Peer-reviewed literature, i.e., journal and conference papers, account for about a third. Most references date from the 2010s, but given that the current decade is only four years old, it is remarkable that almost 40% were published in 2020 or later. A detailed overview of cost data and the corresponding references is given in the Appendix C.
@@
Overnight construction cost (OCC) data are provided on different "readiness levels," i.e., either as so-called "first-of-a-kind" (FOAK) or "n-th-of-a-kind" (NOAK) estimates. FOAK cost estimates acknowledge that the first built reactor of a given type will likely experience unexpected complications and thus be more expensive than subsequent projects of the same design. NOAK cost estimates consequently assume that some learning process reduces unit cost [110]. Learning rates for nuclear vary and, depending on the assessment and data, range from 2 to 15% [111]. However, negative learning rates have also been observed [112]. n indicates the number of units produced to reach a certain cost level. In general, NOAK costs are calculated following Eq. 1 with the cost for the n-th produced unit NOAKn, the cost for the first produced unit FOAK, learning rate x and doubling d of production volume n [110].
@@
Regardless of the readiness level, every value constitutes either an estimation or projection, as none of the discussed reactors have been built. Furthermore, from our gathered literature, we find that only Stewart et al. [94] and Stewart and Shirvan [111] attribute a concrete value to n, i.e., 10. Additionally, in both references, the FOAK starting point is far off from historical precedence, especially for high-capacity LWRs [8]. Regarding learning rates and the feasibility of industry cost reduction projections, refer to Steigerwald et al. [5] who assess learning rates and potential cost reductions for 19 SMR and advanced reactor concepts. For the following, cost data was normalized and adjusted to 2019-USD following the "cost escalation approach" by Abou-Jaoude et al. [113] to ensure comparability.  [^1] If not otherwise indicated, all USD values are in 2019-USD.
Figure 2 shows the variation of cost data between reactor types and readiness levels. FOAK cost data are not available for MSRs. In general, NOAK estimates are more numerous than FOAK estimates. Further, on average, NOAK estimates are lower than FOAK estimates by several thousand USD per kW . This indicates widespread expectations of future cost digressions despite lack of precedence [114, 112, 12]. Figure 3 shows the distribution of OCC values over time. The x-axis shows either the publication year or, if available, the stated reference year. Tables 4, 5 and 6 in Appendix C summarize cost data for OCC, as well as for fixed and variable operational costs for various reactor types. Fewer data points were identified for these cost data than for OCC, and information could not be found for each reactor type.
@@
3 Methodology
The following section describes the methodology used to assess the economic viability of non-light water reactors in a future European energy system. First, we introduce the applied energy system model. Then, we motivate the selection of four reactor technologies for our application based on previous sections, and finally, we describe our cost assumptions, also based on previously discussed analyes.
3.1 Applied Energy System Model
This work applies a comprehensive energy system model to assess different reactor concepts. It utilizes the AnyMOD.jl framework for energy system modeling [22]. Fig. 4 gives an overview of the energy vectors in the model. Sources of primary energy supply in the left column include nuclear and renewables, i.e., solar, wind, hydro, and biomass. The middle column of Fig. 4 lists the secondary energy carriers that can be generated from the primary supply and include electricity, district heat, and various synthetic fuels, including hydrogen, synthetic methane, and methanol. Finally, the right column shows the types of final energy demand that the model must cover. These include final demands for electricity, space heat, and process heat, respectively. The demand for process heat is further subdivided into three levels: low temperature up to 100°C, medium temperature from 100°C to 500°C, and high temperature above 500°C. Finally, the model also covers the demand for transport services, specifically private road, public road, and public rail transport for passengers and freight transport on heavy road, light road, and rail. As indicated in the figure, the model considers nuclear power not just as a source of electricity but also for the provision of district and process heat, as described in Section 3.2. [8]
Figure 4: Energy vectors considered in the applied model.
The model is formulated as a linear optimization problem that decides on capacity investment and operation of capacities to satisfy a fixed final demand at the lowest cost possible. Investment options are different technologies to transport, convert, store, and generate different energy carriers, see Fig. 4. Technology data and demand are based on projections for the year 2040. The Appendix D gives an overview of the considered technologies besides nuclear power.
The spatial scope covers the European Union (excluding isolated island states), the U.K., Switzerland, Norway, and the Balkan region. The temporal scope consists of a single future year in a greenfield setting. Due to their long lifetimes, existing hydropower plants and power grids are available without additional investment. The model uses hourly profiles based on historical data from 2008 for electricity and heat demand as well as renewable capacity factors.
The spatial resolution of the model varies by context. For the characteristics of renewables, all types of heat and hydrogen, the spatial scope is split into 96 clusters. For the electricity balance and transport demand, the scope is split according to the zonal configuration of the European power market. Appendix E details the spatial configuration and the available transport infrastructure to exchange energy between regions. [8, 115]
3.2 Applied Nuclear Technologies and Assumptions
From the analysis of use cases and available technologies (see Sections 2.1 and 2.2) and the available cost data presented in Section 3.3 below, we derive the necessary input parameters for a total of four different reactor technologies to be implemented into the model. These are high-capacity (advanced) light-water reactors (LWR), low-capacity light-water reactors (SMR), sodium-cooled fast reactors (SFR), and high-temperature gas-cooled reactors (HTR). Each technology is assumed to be able to supply a particular form of heat in addition to electricity, see Table 1.
Technology Energy Carrier
LWR Electricity; districtHeat
SMR Electricity; processHeat_low
SFR Electricity; processHeat_medium
HTR Electricity; processHeat_medium; processHeat_high
Table 1: Nuclear technologies included in the energy system model and corresponding energy carrier generation. Abbreviations used: LWR = High-capacity light water reactor; SMR = Small Modular reactor (limited to light water technology); SFR = sodium-cooled fast reactor; HTR = high-temperature reactor
Nuclear power is sometimes portrayed as a dispatchable technology that could bring benefits to an energy system characterized by a high penetration from renewables [19, 116, 117]. Further assumptions relate to individual reactor technologies’ abilities to switch operational modes. We assume that reactors can flexibly switch from electricity to heat provision and vice versa without any downtime or incurred cost. Providing heat increases thermal efficiency levels by an average of 20%. Additionally, the reactors in the model operate without ramp-up time at a universal 90% capacity factor.
3.3 Nuclear Cost Assumptions
@@
From obtained OCC data, total capital costs are calculated following Rothwell [118] by accounting for a uniform seven years of construction time and weighted average cost of capital of 10%. Input parameters that differ depending on reactor type and readiness level are given in Table 2. Due to the heterogeneity of provided data, OM costs are combined from fixed and variable values following the approach used in Göke et al. [8]. They explicitly exclude fuel costs that are included as a spererate value. Further, we assume uniform capacity factors of 90% and operational lifetimes of 40 years. To match other model input cost parameters, cost data for nuclear given in USD-2019 were converted to EUR-2020 with an inflation rate of 1.4% and an exchange rate of 0.8929.
Readiness Level Input Parameter SFR LWR SMR HTR
FOAK OCCa 9,511.47 5,622.9 9,241.42 9,438.52
OMb 50.48 27.03 47.49 45.29
Fuel costb 11.9
NOAK (mean) OCCa 4,677.97 4,983.78 4,407.51 5,649.65
OMb 19.93 20.45 20.82 33.65
Fuel costb 9.16
NOAK (min) OCC 1,476 1,782.62 1,940 2,501.33
OMb 13.31 8.37 15.05 3.13
Fuel costb 9.16
@@
Table 2: Input parameters used for each reactor type and readiness level where differences occur. aGiven in USD per kW. bGiven in USD per MWh. Abbreviations used: LWR = High-capacity light water reactor; SMR = Small Modular Reactor (limited to light water technology); SFR = sodium-cooled fast reactor; HTR = High-temperature reactor
@@ 4 Results
This section presents the results for three overall cost scenarios and model runs. The first assumes mean FOAK costs for each technology, the second mean NOAK costs, and the third minimum NOAK costs, see Table 2. Each of these scenarios contains five model runs. The resolution of each run is 4392 hours. While the first run assumes all four reactor technologies may be built, the other four each exclude one type (e.g., "noHTR" forbids the construction of HTRs). The model runs reflect potential variations in availability and indicate which nuclear technologies are dominated by the others. The results show the strong dependency of system design on input costs. Over all subsequently discussed technology input variations, FOAK cost assumptions lead to a de-facto exclusion of all assumed nuclear power technologies. In contrast, nuclear capacity is added to the system once mean NOAK costs are applied. In the following, we analyze the generation of electricity and heat and necessary capacity developments.
Figure 5 shows the cumulative share of nuclear in total electricity generation, while the composition of the electricity mix is shown in Figure 6. In FOAK scenarios, total electricity generation is around 7,472 TWh. For mean NOAK scenarios, electricity generation is at approximately 6,200 TWh, apart from scenario noHTR in which an additional 581 TWh of electricity is generated to cover the missing high process heat provision from HTRs. The lowest electricity generation of around 5,712 TWh is produced in the noSFR scenario for minimum NOAK cost. In these low-cost scenarios, average electricity generation lies at 5,918 TWh, with one outlier in the noHTR scenario (6,518 TWh).
Analogously, Figures 7 and 8 show the same for heat provision. In all FOAK scenarios, electricity and heat generation from nuclear power is non-existent because no nuclear capacity is built at these cost levels, see Figure 9. Instead, heat demand is covered through a higher degree of power-to-x measures and a higher overall renewable electricity generation. Total heat generation is less variable across all scenarios. On average, 3,559 TWh of heat are generated, with a maximum of 3,589 TWh in the "noLWR" scenario for mean NOAK costs, and a minimum of 3,480 TWh for the "noHTR" scenario at minimum NOAK costs. Notably, LWRs become relevant for district heat generation only when SFRs are excluded, and costs are at minimum NOAK costs, while SMRs remain irrelevant across all scenarios. When nuclear is built, HTRs and SFRs dominate.
Once costs decrease to mean NOAK levels, the model adds nuclear power to the energy mix. However, most of this capacity refers to HTRs. While OCC input parameters for HTRs are higher than for all other reactor technologies, their ability to produce medium- and high-temperature process heat makes them competitive from an energy systems perspective as for these energy carriers, equivalent renewable technologies, i.e., high-temperature heat pumps or hydrogen, are comparatively expensive or also non-existent as of today. 1,109 GW of HTR capacity is built when all technologies are allowed, of which only 49 GW are used primarily for electricity, the remainder for heat provision. This HTR fleet is aided by 63 GW of SFRs that provide mainly medium-temperature process heat (61 GW dedicated to heat). If HTRs are excluded, additional SFR capacity is added (410 GW in total). Still, they cannot fully compensate, resulting in a substantially lower nuclear share in electricity and heat, down from 2% and 30% to 1% and 14%, respectively. LWRs and SMRs play little to no role due to their non-competitiveness with renewables for district and low-temperature process heat provision.
The dominance of HTRs in the installed nuclear fleet ends once costs drop to the minimum NOAK cost assumptions. Here, SFRs become competitive for electricity and heat provision. While they provide only 9% of nuclear heat in the "all" scenario, see Figure 7, they provide the overwhelming majority of nuclear electricity, adding to the substantial share of 23%. If HTRs are excluded, SFRs become even more dominant and are the only nuclear technology to be added to the grid (with a negligible 438 MW of SMRs added). Notably, high-capacity LWRs are introduced for the first time in these scenarios. SMRs are de facto not built in any scenario; they appear to be strictly dominated by other nuclear technologies, and low-temperature heat is provided by cheaper (renewable) alternatives, mainly heat pumps. When SFRs are excluded, the installed nuclear capacity is dominated by HTRs (1335 GW), and 395 GW of LWR capacities. Regardless, HTRs provide most process heat, while LWRs produce some district heat. 5 GW of SMR capacity is built. For all readiness levels, excluding SMR and LWR technologies has no noticable impact.
Figure 5: Share of nuclear power in total electricity production per reactor type and scenario. Abbreviations used: FOAK = "first of a kind"; NOAK = "n-th-of-a-kind"; SFR = sodium-cooled fast reactor; LWR = high-capacity light water reactors; SMR = small modular reactor; HTR = high temperature reactor
Figure 6: Electricity mix for scenarios allowing all reactor technologies ("all") for different readiness levels.
Figure 7: Share of nuclear power in total heat production per reactor type and scenario. Heat includes low, medium and high process heat as well as district heat. Abbreviations used: FOAK = "first of a kind"; NOAK = "n-th-of-a-kind"; SFR = sodium-cooled fast reactor; LWR = high-capacity light water reactors; SMR = small modular reactor; HTR = high temperature reactor
Figure 8: Heat mix for scenarios allowing all reactor technologies ("all") for different readiness levels and types of heat.
Figure 9: Installed nuclear capacity used for electricity and heat for scenarios and readiness levels compared to currently installed capacity in Europe as of April 2024 [119]. Note that FOAK capacities have been excluded from this figure because no nuclear is built. Abbreviations used: NOAK = "n-th-of-a-kind"; SFR = sodium-cooled fast reactor; LWR = high-capacity light water reactors; SMR = small modular reactor; HTR = high temperature reactor.
Total system costs for the FOAK scenarios are the highest, averaging at €317.5 billion with a standard deviation (sd) of €21.1 million. Mean NOAK scenarios average €307.2 billion (sd = €2.7 billion), and minimum NOAK scenarios are the cheapest, averaging only €290 billion (sd = €3.1 billion). The main differences in cost stem from the increased expansion of nuclear power capacity at comparatively low cost (compared to actual to be expected costs, see Section 3.3). In the FOAK scenarios, additional flexibility measures are built, such as open-circuit hydrogen power plants, and consequently, investments into powerToX technologies become necessary. Gas power plants are de facto not built in the minimum NOAK scenarios.

### v2

@@ 2.2.2 Reactors with Fast Neutron Spectra
Lead-cooled fast reactors (LFR) operate similarly to SFRs, but use lead as a coolant. Lead is less reactive than sodium and could thus reduce operational risks and necessary safety barriers [78, 25]. However, operational experience is almost exclusively limited to Russia, where LFRs were experimented with for submarine propulsion, and a single reactor is currently under construction [74]. Further, a demonstration LFR with an accelerator-driven system is currently under development in Belgium (MYRRHA) [77]. Due to limited operational experience, cost information is scarce. However, it is expected that LFRs could be cheaper than SFRs due to reduced safety risks, and if a closed fuel cycle can be implemented, LFRs might also compete with current LWRs [78].
@@ 2.2.4 Molten salt reactors
Molten salt reactors (MSR) use molten salt as a coolant [82]. MSRs were initially developed for aircraft propulsion, and concepts vary regarding moderator, fuel, neutron spectrum, and fissile material. There are no commercially operational MSRs, but several concepts are currently under development in China, Switzerland, and Russia, amongst others [77, 83]. MSRs could become applicable for heat provision in decarbonized energy systems as the proposed salts melt at around 500°C and boil at around 1400°C, allowing for outlet temperatures ranging from 600 to 700°C [82, 77]. Most MSR concepts are still in the early development stages [83], so information on their economic competitiveness is limited. While in the 1970s, capital costs of MSR were estimated in the same range as LWRs [84], current concepts are estimated to be cheaper. For example, the so-called "Advanced High-Temperature Reactor," also referred to as "Fluoride High-Temperature Reactor," is estimated to cost less than 1000 USD per kW, and the ThorCon concept MSR, developed by company Martingale, is envisioned to produce electricity for as little as 30 to 50 USD per MWh [85, 77].
3 Methodology
The following section describes the methodology used to assess the potential of non-LWR concepts in a future European energy system. First, we introduce the applied energy system model. Then, we motivate the selection of four reactor technologies for our application based on previous sections, and finally, we describe our cost assumptions based on our vast cost analysis.
3.1 Applied Energy System Model
This work applies a comprehensive energy system model to assess different reactor concepts [86]. It utilizes the AnyMOD.jl framework for energy system modeling [31]. Fig. 2 gives an overview of the energy vectors in the model. Sources of primary energy supply in the left column include nuclear and renewables, i.e., solar, wind, hydro, and biomass. The middle column of Fig. 2 lists the secondary energy carriers that can be generated from the primary supply and include electricity, district heat, and various synthetic fuels, including hydrogen, synthetic methane, and methanol. Finally, the right column shows the types of final energy demand that the model must cover. These include final demands for electricity, space heat, and process heat, respectively. The demand for process heat is further subdivided into three levels: low temperature up to 100°C, medium temperature from 100°C to 500°C, and high temperature above 500°C. In addition, the model also covers the demand for transport services, specifically private road, public road, and public rail transport for passengers and freight transport on heavy road, light road, and rail. Finally, it includes an exogenous fuel demand covering aviation, navigation, and feedstocks in the industry. As indicated in the figure, the model considers nuclear power not just as a source of electricity but also for the provision of district and process heat, as described in Section 3.2.
Figure 2: Energy vectors considered in the applied model.
The model is formulated as a linear optimization problem that decides on capacity investment and operation of capacities to satisfy a fixed final demand at the lowest cost possible. Investment options are different technologies to transport, convert, store, and generate different energy carriers, see Fig. 2. Technology data and demand are based on projections for the year 2040. Appendix E gives an overview of the considered technologies besides nuclear.
The model pursues a stochastic approach to capture the effect of uncertain weather conditions on renewable generation and demand based on Göke et al. [86]. Weather foresight is limited to a month, and the model explicitly models 16 different months to account for extreme seasonal variations. However, a robust formulation ensures reliability for all possible combinations of these months, resulting in 12 different climate years. The model employs time-series clustering to reduce the underlying time series from 8760 to 2856 hours, while maintaining daily patterns.
The spatial scope covers the European Union (excluding isolated island states), the U.K., Switzerland, Norway, and the Balkan region. The temporal scope consists of a single future year in a greenfield setting. Due to their long lifetimes, existing hydropower plants and power grids are available without additional investment.
The spatial resolution of the model varies by context. For the characteristics of renewables, all types of heat and hydrogen, the spatial scope is split into 96 clusters. For the electricity balance and transport demand, the scope is split according to the zonal configuration of the European power market. Appendix F details the spatial configuration and the available transport infrastructure to exchange energy between regions. [9, 87]
3.2 Applied Nuclear Technologies and Technical Assumptions
From the analysis of use cases and available technologies (see Sections 2.1 and 2.2) and the available cost data presented in Section 3.3.2 below, we derive the necessary input parameters for a total of four different reactor technologies to be implemented into the model. These are high-capacity (advanced) light-water reactors (LWR), low-capacity light-water reactors (SMR), sodium-cooled fast reactors (SFR), and high-temperature gas-cooled reactors (HTR). Each technology is assumed to be able to supply a particular form of heat in addition to electricity, see Table 1.
Nuclear power is sometimes portrayed as a dispatchable technology that could bring benefits to an energy system characterized by a high penetration from renewables [22, 88, 89]. Further assumptions relate to individual reactor technologies’ abilities to switch operational modes. We assume that reactors can flexibly switch from electricity to heat provision and vice versa without any downtime or incurred cost. We allow all plants apart from high-capacity LWRs to operate in three distinct modes that influence the thermal efficiency of the reactors. A full heat mode increases thermal efficiency by an average of 26% compared to the mode with the highest output of electricity. A medium mode increases thermal efficiencies by 16%. LWRs may operate only on the medium and low heat modes. The reactors in the model operate without ramp-up time at a universal 90% capacity factor independent of the chosen mode.
3.3 Nuclear Cost Data
3.3.1 Cost Data Analysis
@@
Over the last decade, nuclear power has gained momentum in energy models as a low-carbon technology [8]. Many of these models assume substantial cost reductions for nuclear in the coming decades down to overnight costs of 5,000 or even 2,000 USD per kW (e.g., Bistline et al. [23]) and highlight the necessity of cost reductions for nuclear to be competitive in energy systems with high shares of renewables [2, 90].
@@
Thus, to assess the economic efficiency of non-light water reactors in the energy system model (see Section 3.1), it is first necessary to determine relevant cost data and other relevant information on potentially viable reactor types. Consequently, we conducted a literature analysis of cost projections and assumptions for various reactor types. Previous research has shown that for high-capacity LWRs, cost projections and experience differ strongly [9]. For SMR concepts, Steigerwald et al. [6] and Kim and Macfarlane [17] show that to-be-expected costs and industry expectations also vary substantially. This is in line with findings by Abdulla et al. [91] who determine that cost projections for nuclear are frequently overly optimistic and diverge strongly.
@@
The here presented cost analysis is based on 65 references [92, 4, 93, 94, 95, 96, 97, 80, 98, 99, 100, 101, 102, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 18, 119, 120, 121, 122, 123, 124, 125, 126, 127, 128, 129, 130, 131, 132, 81, 16, 133, 134, 135, 136, 77, 137, 138, 139, 140, 61, 91, 141, 142, 143, 144, 25, 145, 146, 26]. About 51% are so-called grey literature, mostly industry and research organization reports. Peer-reviewed literature, i.e., journal and conference papers, account for about a third. Almost 45% of the references are only five years old. A detailed overview of cost data and the corresponding references is given in Appendix C.
@@
Overnight construction cost (OCC) data are provided on different "readiness levels," i.e., either as so-called "first-of-a-kind" (FOAK) or "n-th-of-a-kind" (NOAK) estimates. FOAK cost estimates acknowledge that the first built reactor of a given type will likely experience unexpected complications and thus be more expensive than subsequent projects of the same design. NOAK cost estimates consequently assume that learning processes reduce unit costs [147]. Learning rates for nuclear vary and, depending on the assessment and data, range from 2 to 15% [111]. However, negative learning rates have also been observed [148]. n indicates the number of units produced to reach a certain cost level. In general, NOAK costs are calculated following Eq. 1 with the cost for the n-th produced unit NOAKn, the cost for the first produced unit FOAK, learning rate x and doubling d of production volume n [147].
@@
Regardless of the readiness level, every value constitutes either an estimation or projection, as none of the discussed reactors have been built. Furthermore, from our gathered literature, we find that only Stewart et al. [119] and Stewart and Shirvan [111] attribute a concrete value to n, i.e., 10. Nøland et al. [143] investigate the required number of built reactors for four types of SMRs. They find that while only eight units of the Rolls-Royce SMR (470 MWe) would suffice to reach announced cost projections, the same approach would require 41,025 VOYGR units. Additionally, in all three references, the FOAK starting point is far off from historical precedence, especially for high-capacity LWRs [9]. Regarding learning rates and the feasibility of industry cost reduction projections, refer to Steigerwald et al. [6] who assess learning rates and potential cost reductions for 19 SMR and advanced reactor concepts. For the following, cost data was normalized and adjusted to 2019-USD following the "cost escalation approach" by Abou-Jaoude et al. [149] to ensure comparability.  [^1] If not otherwise indicated, all USD values are in 2019-USD.
Figure 3 shows the variation of cost data between reactor types and readiness levels. FOAK cost data are not available for MSRs. In general, NOAK estimates are more numerous than FOAK estimates. Further, on average, NOAK estimates are lower than FOAK estimates by several thousand USD per kW. This indicates widespread expectations of future cost digressions despite lack of precedence [150, 148, 13]. Figure 4 shows the distribution of OCC values over time. The x-axis shows either the publication year or, if available, the stated reference year. Tables 3, 4 and 5 in Appendix C summarize cost data for OCC, as well as for fixed and variable operational costs for various reactor types. Fewer data points were identified for these cost data than for OCC, and information could not be found for each reactor type.
@@
3.3.2 Nuclear Cost Assumptions
@@
From obtained OCC data, total capital costs are calculated following Rothwell [151] by accounting for a uniform seven years of construction time and weighted average cost of capital of 10%. We conduct a step-by-step scenario approach to determine cut-in costs for different nuclear technologies. In ten deterministic scenarios, we reduce OCC for each considered nuclear technology from the mean FOAK value to the minimum NOAK via linear interpolation. Further relevant input parameters are given in Table 1 and reflect mean NOAK values from the literature. Due to the heterogeneity of provided data, OM costs are combined from fixed and variable values following the approach used in Göke et al. [9]. These OM costs explicitly exclude fuel costs that are included as a separate value. Further, we assume uniform capacity factors of 90% and operational lifetimes of 40 years. To match other model input cost parameters, cost data for nuclear given in USD-2019 were converted to EUR-2020 with an inflation rate of 1.4% and an exchange rate of 0.8929.
Reactor Technology OCCa O&M costb Fuel Costb Energy Carriers
LWR 1747 - 5511 18.51 8.29 Electricity & District Heat
SMR 1902 - 9126 18.85 Electricity & Low-temperature process heat
SFR 1446 - 9323 18.05 Electricity & Medium-temperature process heat
HTR 2451 - 9252 30.47 Electricity & High-temperature process heat
@@
Table 1: Cost-related input parameters used for each reactor type. aGiven in USD per kW. bGiven in USD per MWh. Abbreviations used: LWR = High-capacity light water reactor; SMR = Small Modular Reactor (limited to light water technology); SFR = sodium-cooled fast reactor; HTR = High-temperature reactor
@@ 4 Results
In the following section, we first present the results of the model analysis for nuclear capacity expansions before discussing the impact of nuclear expansion on the whole energy system.
4.1 Overview
Figure 5: Shares of nuclear technologies in electricity and heat provision plotted against assumed OCC values for ten cost scenarios.
Overall, we find that with decreasing costs for nuclear, its shares in both electricity and heat provision increase. Figure 5 shows the share of nuclear, i.e., the combined production of all four technologies, for electricity, the combined provision of heat, and all four respective types of heat. It also shows the corresponding stepwise digression of OCC values for each technology.
The share of electricity production from nuclear reaches the current levels in the EU (22.6% as of 2023 [32]) once OCC values are almost at the NOAK level (No. 2 in Figure 5). In this scenario, OCC values are 2,165 USD/kW for LWR, 3,207 USD/kW for HTR, 2,704 USD/kW for SMR, and 2,322 USD/kW for SFR.
In contrast, the nuclear share of heat generation rises continuously and reaches its maximum level of 43.2%, again when costs are almost at the NOAK level. High and medium process heat, provided by HTRs and SFRs, respectively, reach high shares at higher costs (No. 6 & 7).
The share of nuclear in low temperature process heat provision remains marginal until costs for SMR fall below 2,704 USD/kW, when the share is determined at 3.8% (No. 2). The highest share is 52.9% at the lowest costs of 1,902 USD/kW (No. 1).
It is noteworthy that the nuclear heat share falls between the cheapest (No.1) and second-to-last cost scenario (No. 2). This is driven by the substantial reduction of district heat provision from LWRs from a maximum of 18% to about 0.43% in the lowest cost scenario. As shown in Figure 5, the OCC for LWRs supersede those of SFRs in the minimum NOAK scenario (No. 1). Consequently, the cost-optimizing model selects the superior technology, SFR, to provide the electricity that was previously provided by LWRs. Given the relatively low share of district heat provision by LWRs in other scenarios, and the model’s switch to another reactor technology, the main purpose of LWRs was to generate electricity. Until the minimum NOAK cost scenario, LWRs remained as the cheapest reactor technology. The production of district heat seems to have been of secondary priority. As can be seen in Figure 11 in Appendix D, competitive technologies for district heating compensate for the reduction of LWR capacities in the lowest cost scenario (No. 1).
Low-temperature process heat provision remains marginal throughout the scenarios until the lowest cost scenario when SMRs become cheaper than competitive technologies (No. 1). Only then are measurable capacities built. Figure 6 shows the nuclear capacities for the different scenarios. The dominance of LWRs for electricity generation is prominent until the lowest cost scenario, when SFRs become cheaper than LWRs (No. 1). Total installed nuclear capacities reach currently installed European capacities (EU and UK, i.e., 102.2 GW [32]) between scenarios No. 6 & 7. LWRs are built once their OCC fall enough to compete with alternative electricity generation technologies, i.e., around 3,000 USD/kW (No. 5). But, as mentioned above, LWRs are no longer built once SFRs become even cheaper. In the lowest cost scenario, total nuclear capacities supersede 1,523 GW; driven mainly by very low cost assumptions for SFRs that produce mainly electricity.
As shown in Figure 5, medium and high process heat provision is swiftly saturated which corresponds to the maximum capacity of HTRs of around 27 GW being reached fairly early (No. 5), and SFR capacity increasing only once the technology becomes cheaper than LWRs (No. 1) (see also Figure 6).
Figure 6: Installed nuclear capacities for the different cost scenarios.
4.2 System Impact of Nuclear
Total system costs for the scenarios with low nuclear shares are the highest. However, until installed nuclear capacity reaches levels above 200 GW, these costs remain at an average of €303.6 billion with a standard deviation (sd) of €1.9 billion. In the lowest cost scenario, overall system costs lie at €264.8 billion. Nuclear capacity expansions occur at the expense of open-cycle hydrogen power plants, other gas power plants, solar and wind capacities as well as utility-operated heat plants. Cross-border exchange of electricity is reduced the more nuclear is built. Fuel import dependencies are also reduced, although they remain at similar levels for the last five scenarios.
Lower system costs are to some extent reflected by the reduced necessity of providing system flexibility for intermittent renewables, such as solar and wind. We assume favorable operational parameters for nuclear, such as a capacity factor of 90%, that allows the system to run nuclear capacities at almost full capacity. This is shown in Figure 7(a) that depicts a week in February for the German grid in the lowest cost scenario (No. 1). Compared to Figure 7(b), showing the results for the highest cost scenario (No. 10) and no nuclear in the system, one can see how nuclear blocks the grid instead of complementing intermittent technologies.
A similar pattern occurs for the different types of heat provision. Heat and electricity, however, are not directly comparable due to inertia, reflected in the model via four-hour instead of hourly timesteps. Regardless, the dominance of nuclear, once built, in heat provision can be observed in Figures 8(a), 9(b), 10(b), and 11(b). We show the highest and lowest cost scenarios for each type of heat provision but for simplicity, we only show high process heat generation in this section. Refer to Appendix D for figures on low and medium process heat as well as district heat. We find that nuclear does not take advantage of the granted flexibility and blocks the grid. This is in line with findings of previous work [9].
(a) Least cost scenario (minimum NOAK, No. 1)
Figure 7: Exemplary electricity generation in Germany
(b) Highest cost scenario (mean FOAK, No. 10)
(a) Least cost scenario (minimum NOAK, No. 1)
Figure 8: Exemplary high-temperature process heat generation in Germany
(b) Highest cost scenario (mean FOAK, No. 10)